Find polynomials q and r such that f(x)=q(x)g(x)+r(x) where the degree of r is strictly less than degree of g.
f(x)=x^4-3x^3+2x^2-x+1 and g(x)=x^2-3

Answer by stanbon(75887)

Find polynomials q and r such that f(x)=q(x)g(x)+r(x) where the degree of r is strictly less than degree of g.
f(x)=x^4-3x^3+2x^2-x+1 and g(x)=x^2-3
---
Divid
f(x) by g(x) to find the quotient(q(x)) and the remainder(r(x)).
----
Ans:
q(x) = x^2 - 3x +8
r(x) = -x + 25
------
Cheers,
Stan H.
------------